At 6 A.M. a snowplow, traveling at a constant speed, begins to clear a highway leading out of town. At 8 A.M. an automobile begins traveling the highway at a speed of 35 mi/hr and reaches the plow 20 minutes later. Find the speed of the snowplow.
Answer by richwmiller(17219)

6am
8am - 8:20 20/60=1/3
snow plow r*t
35 mph
35*1/3=11 2/3 miles
r*t=d
plow x mph for 2 hours hours went y miles
car 35 mph 35-p=g miles gaining
y/g=1/3 hours
r*(7/3)=35/3
r*7=35
r=5 mph for snow plow
(r+g)*1/3=11 2/3
35/3=11 2/3
r+g=35
5+g=35
g=30
30/3=10 miles ahead check 5*2=10
distance was 11/2/3 m
snow plow went 5 mph